solve the equestion using the addition property of equality.be sure to check your proped solution.
-4y+8=-5y+1

Answer by Maths68(1474)

-4y+8=-5y+1
Add -8 both sides
-4y+8+(-8)=-5y+1+(-8)
-4y+8-8=-5y+1-8
cancel out +8 and -8 on RHS
-4y=-5y-7
Add 5y both sides
-4y+5y=-5y-7+5y
cancel out +5y and -5y on LHS
y=-7



Check
======
-4y+8=-5y+1
Put y=-7
-4(-7)+8=-5(-7)+1
28+8=35+1
36=36